# Achieving Optimal Sample Complexity for a Broader Class of Signals in Sparse Phase Retrieval

### v2

@@ VI-A Experimental Setup
Across all simulations, the signal dimension is fixed at $n=1000$. The measurement vectors $\{\mathbf{a}_{i}\}_{i=1}^{m}\subset\mathbb{C}^{n}$ are i.i.d. standard complex Gaussian vectors. The $k$-sparse signal $\mathbf{x}\in\mathbb{C}^{n}$ is constructed by first selecting a support of size $k$ uniformly at random. The non-zero entries are then generated from one of three models: i) i.i.d. standard complex Gaussian, ii) all ones (binary), and iii) exponentially decaying magnitudes. All reported results are averaged over $1000$ independent trials to mitigate random fluctuations.
We compare gESP with CoPRAM, TPM, and ESP. For a fair comparison, we consider only the initialization stage of CoPRAM, and use the recommended hyperparameters for TPM. For gESP, we set the theoretically optimal parameter $p_{opt}$ from (38) for each trail.
@@
$\text{Relative Error}=\frac{\|\mathbf{z}-\mathbf{x}\|_{2}}{\|\mathbf{x}\|_{2}},$
@@ VI-B Performance Comparison
In this experiment, we vary the sampling ratio $m/n$ from $0.05$ to $1.0$ and record the corresponding relative error.
@@
As depicted in Figs. 3a3c, the performance on sparse Gaussian signals provides a direct and strong validation of our theoretical analysis. For this signal class, gESP was configured using the parameter $p$ determined by our theoretically derived optimum. Consistent with the predictions of our theory, gESP achieves the best performance, consistently and slightly outperforming ESP and other state-of-the-art methods. This clear correspondence between our theory and the empirical results underscores the accuracy of our analysis for signals with rich amplitude variations.
@@
In contrast, the results for binary and exponentially decaying signals, while not perfectly aligned with asymptotic predictions, offer valuable insights into the practical behavior of the algorithms. For binary signals, theory suggests the complexity would always be $\Omega(k^{2}\log n)$ no matter what value $p$ is, and thus one might expect similar performance across methods. However, we set $p=k$ in gESP, and it demonstrates a surprisingly large performance margin. We conjecture that this significant practical advantage stems from the influence of non-asymptotic constant factors that our analysis does not model but which clearly favor the gESP framework. Similarly, for exponentially decaying signals, gESP is marginally outperformed by ESP. This is also likely an effect of such constants; our theory provides the optimal scaling for $p$, but in this finite-dimensional regime, the true optimal value is empirically found to be $p=1$ (i.e., ESP). This does not contradict our theory but rather highlights that while it provides a powerful general guideline, the precise optimal parameter in practice can be influenced by the signal’s specific structure.
@@ VII Conclusion
In this paper, we investigate the problem of sparse phase retrieval, focusing on improving initialization algorithms for recovering a $k$-sparse $n$-dimensional signal from $m$ phaseless observations. Through an analysis of existing methods, we propose generalized Exponential Spectral Pursuit (gESP), which enhances initialization performance and refines the understanding of sample complexity in sparse phase retrieval. Empirical evaluations further validate the practical effectiveness of gESP. Simulation experiments demonstrated its robustness and efficiency in recovering sparse signals, highlighting its potential utility in real-world applications.
While this work advances both theoretical bounds and practical performance for sparse phase retrieval, it does not fully resolve the statistical-to-computational gap. Specifically, under Gaussian measurements, $\Omega(k\log n)$ samples are theoretically sufficient to determine the target signal. However, in practical scenarios requiring polynomial-time algorithms, the sample complexity increases to $\Omega(k^{2}\log n)$ in the worst case, such as when the target signal consists of 0-11 entries. Future research could aim to bridge this gap by exploring novel algorithmic frameworks or leveraging insights from other areas, such as computational complexity and optimization. Additionally, extending the applicability of gESP to broader measurement models and refining the refinement stage offer promising directions for advancing this field.
@@ Appendix A Basic tools and Lemmas
In this section, we present several useful lemmas and necessary proofs. First of all, recall that
@@
where $\lambda^{2}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$. It is challenging to derive the expectation of $\mathbf{Z}$. Following the technique in [12, 28], we introduce
@@
In the following, we introduce several results concentrating on the effect of the ceiling sign on the sample complexity.
@@
$\Omega(s(\lceil\sqrt{k}\rceil)$ $=\Omega(s(c\lceil\sqrt{k}\rceil)),$ (A.6)
@@
Note that $\sqrt{k}\leq\lceil\sqrt{k}\rceil\leq\sqrt{k}+1$, then (A.5) can be easily derived from the following relationship.
@@
Since $x_{(j)}$’s are the rearrangement of the entries of $\mathbf{x}$ with a descend order in their modulus, the summation in the denominator can be estimated as
@@
In that $c$ is a constant as $k$ and $n$ tend to infinity, (A.6) is a natural derivation. ∎
Lemma 2 (Bernstein inequality for complex cases).
Let $X_{i}$ be independent, mean-zero complex sub-exponential random variables for $i\in[m]$, then for any $\epsilon>0$, we have
$\mathbb{P}\left(\left\lvert\frac{1}{m}\sum_{i=1}^{m}X_{i}\right\rvert\geq\epsilon\right)\leq 4\exp\left(-cm\min\left\{\frac{\epsilon^{2}}{K^{2}},\frac{\epsilon}{K}\right\}\right),$
where $K=\mathop{\max}_{i\in[m]}\|X_{i}\|_{\psi_{1}}$ and $c$ is the numerical constant. Also see [22, Lemma 10].
Next, the difference between $\lambda^{2}$ and $\|\mathbf{x}\|^{2}$ is analyzed.
@@
Given $k$-sparse signal $\mathbf{x}$, for any constant $\beta>0$, suppose that $\mathbf{a}_{i}$’s are complex Gaussian random variables. Then provided that $m\geq C_{\beta}\kappa^{2}$,
$\left(1-\frac{\beta}{\kappa}\right)\|\mathbf{x}\|^{2}\leq\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}\leq\left(1+\frac{\beta}{\kappa}\right)\|\mathbf{x}\|^{2}$
holds with probability at least $1-e^{-c_{\beta}m}$, where $\kappa$ can be substitute by any quantity satisfying $\frac{\beta}{\kappa}<1$, such as $ps(p)$, and $c_{\beta}$ and $C_{\beta}$ are numerical constants determined by $\beta$.
@@
It is easy to find that $|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}$’s are independent sub-exponential random variables with mean $\mathbb{E}[|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}]=\|\mathbf{x}\|^{2}$ and $c_{0}\|\mathbf{x}\|^{2}$ as their squared sub-exponential norms. Hence, employing Bernstein’s inequality yields
$\mathbb{P}\left(\left\lvert\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}-\|\mathbf{x}\|^{2}\right\rvert>\frac{\beta}{\kappa}\|\mathbf{x}\|^{2}\right)$
$~~~~~~~\leq 2\exp\left(-C_{\beta}m\min\left\{\frac{\beta^{2}}{\kappa^{2}K^{2}},\frac{\beta}{\kappa K}\right\}\right),$
where $K=\mathop{\max}_{i\in[m]}\||\mathbf{a}_{i}^{*}\mathbf{x}|^{2}\|_{\psi_{1}}/\|\mathbf{x}\|^{2}$ and $C_{\beta}$ is the numerical constant. Thus we obtain the desired result.∎
Finally, we introduce a useful lemma that measures the similarity between the exponential spectral $\mathbf{Z}$ and its estimator $\mathbb{E}[\tilde{\mathbf{Z}}$].
@@
Suppose a constant $\eta>0$ and a positive integer $q\leq k$. For all $S\subseteq[n]$ satisfying $|S|=q$, let $T$ be a subset of $\mathbb{C}^{n}$ such that $\|\mathbf{z}\|\leq 1$ holds for all $\mathbf{z}\in T$. Given any constant $\epsilon\in(0,1)$, let $\mathcal{N}(T,\epsilon)$ be the covering number [31, Definition 4.2.2] of $T$ and $g(k,s(p))$ be a function with respect to $k$ and $s(p)$ satisfying $g(k,s(p))\geq 1$. It holds with probability at least $1-2e^{-c_{\eta}m}$ that
$\sup_{\mathbf{z}_{1},\mathbf{z}_{2}\in T}\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{2}\leq\frac{\eta}{g(k,s(p))}$ (A.7)
with the condition that
$m\geq C_{\eta}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)),$ (A.8)
where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
In particular, when $T=\{\mathbf{u}|\operatorname{supp}(\mathbf{u})=S,\|\mathbf{u}\|\leq 1\}$, the inequality
$\left\lVert\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right\rVert\leq\frac{\eta}{2g(k,s(p))}$ (A.9)
holds true given $m\geq C_{\eta}qg^{2}(k,s(p))\log n$.

### v3

@@ VI-A Experimental Setup
In all simulations, the signal dimension is fixed at $n=1000$. The measurement vectors $\{\mathbf{a}_{i}\}_{i=1}^{m}\subset\mathbb{C}^{n}$ are i.i.d. standard complex Gaussian vectors. The $k$-sparse signal $\mathbf{x}\in\mathbb{C}^{n}$ is constructed by first selecting a support of size $k$ uniformly at random. The non-zero entries are then generated from one of three models: i) i.i.d. standard complex Gaussian, ii) all ones (binary on support), and iii) exponentially decaying magnitudes. All reported results are averaged over $1000$ independent trials to mitigate random fluctuations.
We compare gESP with CoPRAM, TPM, and ESP. For a fair comparison, we focus on the initialization stage of CoPRAM, and use the recommended hyperparameters for TPM. For gESP, we set the theoretically optimal parameter $p_{opt}$ from (38) for each trial.
@@
$\text{Relative Error}=\frac{\operatorname*{dist}(\mathbf{z},\mathbf{x})}{\|\mathbf{x}\|_{2}},$
@@ VI-B Performance Comparison
In this experiment, we vary the sample ratio $m/n$ from $0.05$ to $1.0$ and record the corresponding relative error.
@@
As depicted in Figs. 3a3c, the performance on sparse Gaussian signals provides a direct and strong validation of our theoretical analysis. For this signal class, gESP was configured using the parameter $p$ determined by our theoretically derived optimum. Consistent with the predictions of our theory, gESP achieves the best performance, consistently outperforming ESP and other state-of-the-art methods, albeit slightly. This clear correspondence between our theory and the empirical results underscores the accuracy of our analysis for signals with rich amplitude variations.
@@
In contrast, the results for binary and exponentially decaying signals, while not perfectly aligned with asymptotic predictions, offer valuable insights into the practical behavior of the algorithms. For binary signals, the theory suggests that the complexity remains $\Omega(k^{2}\log n)$ regardless of the value of $p$, leading one to expect similar performance across methods. However, we set $p=k$ in gESP, and it demonstrates a surprisingly large performance margin. We attribute this significant practical advantage to the influence of nonasymptotic constant factors that our analysis does not model but which clearly favor the gESP framework. Similarly, for exponentially decaying signals, gESP is marginally outperformed by ESP. This is also likely an effect of such constants; our theory provides the optimal scaling for $p$, but in this finite-dimensional regime, the true optimal value is empirically found to be $p=1$ (i.e., ESP). This does not contradict our theory but rather highlights that while it provides a powerful general guideline, the precise optimal parameter in practice can be influenced by the signal’s specific structure.
VI-C Robustness to the Choice of Parameter $p$
A core feature of gESP lies in that it uses the tuning parameter $p$ to balances the captured signal energy against the accumulated perturbation. In this subsection, we experimentally investigate the sensitivity of gESP to the choice of $p$ and demonstrate its practical performance.
We consider a power-law signal with decay rate $\alpha=0.5$ (i.e., $|x_{(i)}|\propto i^{-0.5}$ for $i=1,\ldots,k$). This structure of signal is particularly selected as it is a representative “heavy-tailed” case: neither essentially single-spiked nor perfectly flat. Such a signal profile effectively challenges the algorithm to balance energy and perturbation, making it an ideal candidate to evaluate the impact of different selection strategies. In the following, we test the performance of gESP with different choices of $p$ on this power-law signal:
  1. 1.
    The standard single-peak approach $p=1$ (corresponding to ESP).
  2. 2.
    The theoretical oracle $p$ obtained by minimizing the bound in Theorem 1.
  3. 3.
    The heuristic surrogate $p=\lceil\sqrt{k}\rceil$ as a middle-ground choice.
  4. 4.
    The full support selection $p=k$.
Figure 4: Performance comparison of gESP with different choices of $p$ under the designed power-law signal. We set $n=1000,k=100$ with $m/n$ varying from $0.05$ to $1.0$. The multi-peak strategies ($p>1$) significantly outperform the single-peak baseline. Notably, the algorithm is robust to large $p$, with the full support choice ($p=k$) achieving the best empirical performance.
The results are shown in Fig. 4. We observe that all multi-peak strategies ($p>1$) significantly outperform the baseline ESP ($p=1$). Notably, the empirical performance remains stable or even improves as $p$ increases, with the full support choice ($p=k$) achieving the best performance in this setting.
This discrepancy between the theoretical oracle and the empirical optimum can be attributed to two main factors. First, our theoretical analysis is asymptotic in nature (assuming $m,n,k\to\infty$), whereas the experiments are conducted in a finite-dimensional setting ($n=1000,k=100$). Clearly the finite-sample behavior may arguably tolerate more perturbation than the asymptotic worst-case prediction. Second, the “oracle” choice is derived by minimizing a theoretical upper bound on the sample complexity. This bound relies on standard concentration inequalities (e.g., Bernstein-type bounds), which are sufficient yet inherently conservative. As a result, the theoretical derivation tends to overestimate the perturbation accumulation associated with large $p$, leading to a more conservative recommendation compared to the algorithm’s actual robustness behavior in practice.
Practically, this result is encouraging: it implies that users can simply select a sufficiently large $p$ (e.g., above the expected sparsity level) to achieve the near-optimal performance without precise knowledge of the energy profile $s(p)$.
@@ VII Conclusion
In this paper, we have addressed the problem of sparse phase retrieval, focusing on improving initialization algorithms for recovering a $k$-sparse $n$-dimensional signal from $m$ phaseless observations. Motivated by the limitations of existing methods that rely on specific structural assumptions of signals, we have proposed generalized Exponential Spectral Pursuit, which promotes the initialization performance and also enhances the understanding of sample complexity in sparse phase retrieval. Empirical evaluations further validate the practical effectiveness of gESP. Simulation experiments demonstrate its robustness and efficiency in recovering sparse signals, highlighting its potential utility in real-world applications.
Although this work advances both theoretical bounds and practical performance for sparse phase retrieval, it does not fully resolve the statistical-to-computational gap. Specifically, under Gaussian measurements, an information-theoretic bound of $\Omega(k\log n)$ samples is sufficient to uniquely determine the target signal. However, for polynomial-time recovery, the sample complexity increases to $\Omega(k^{2}\log n)$ in the worst case, for example, when the target signal consists of binary entries. Future research could aim to bridge this gap by exploring novel algorithmic frameworks or leveraging insights from other areas, such as computational complexity and optimization. Furthermore, extending the applicability of gESP to broader measurement models and optimizing the subsequent refinement stage offer promising directions to advance this field.
@@ Appendix A Basic Tools and Lemmas
In this section, we present several useful lemmas and necessary proofs. Recall the definition of $\mathbf{Z}$:
@@
where $\lambda^{2}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$. Deriving the exact expectation of $\mathbf{Z}$ is challenging due to the dependency between $y_{i}$ and $\lambda^{2}$. Following the technique in [31, 33], we introduce a proxy matrix:
@@
In the following, we introduce several results concerning the effect of the ceiling function on the sample complexity.
@@
$\Omega(s(\lceil\sqrt{k}\rceil))$ $=\Omega(s(c\lceil\sqrt{k}\rceil)),$ (A.6)
@@
Note that $\sqrt{k}\leq\lceil\sqrt{k}\rceil\leq\sqrt{k}+1$, then (A.5) can be easily derived from the limit:
@@
Since $x_{(j)}$’s are the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes, the summation in the denominator can be estimated as
@@
Since $c$ is a constant as $k$ and $n$ tend to infinity, (A.6) follows naturally. ∎
The next lemma gives the error bound of $\lambda^{2}$.
@@
Consider the problem (1). With probability at least $1-n^{-c}$, it holds that
$\left|\lambda^{2}-\|\mathbf{x}\|^{2}\right|\leq C\sqrt{\frac{\log n}{m}}\|\mathbf{x}\|^{2}.$
@@
Let $y_{i}=\mathbf{a}_{i}^{*}\mathbf{x}$. Since $\mathbf{a}_{i}\sim\mathcal{CN}(0,\mathbf{I})$, the variable $y_{i}\sim\mathcal{CN}(0,\|\mathbf{x}\|^{2})$. The term $|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}/\|\mathbf{x}\|^{2}$ follows a standard exponential distribution, which is sub-exponential with constant norm.
We apply the standard Bernstein inequality for the sum of independent sub-exponential random variables [23, Theorem 2.8.1]. For any $t\geq 0$,
$\mathbb{P}\left(\left|\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}-\|\mathbf{x}\|^{2}\right|\geq t\|\mathbf{x}\|^{2}\right)$
$~~~~~~~~~~\leq 2\exp\left(-cm\min(t^{2},t)\right).$
Setting $t=C\sqrt{\frac{\log n}{m}}$ yields the desired result. ∎
The following lemma gives the operator norm bound of the weighted average of random matrices, which is a natural extension of the classical result in random matrix theory.
Lemma 3 (Uniform Concentration with Bounded Weights).
Let $\mathbf{a}_{i}\sim\mathcal{CN}(0,\mathbf{I}_{n})$ be independent standard complex Gaussian vectors. Let $\xi_{i}\in\mathbb{R}$ be independent scalar random variables (potentially depending on $\mathbf{a}_{i}$) satisfying $|\xi_{i}|\leq B$ almost surely for some constant $B>0$.
With probability at least $1-2\exp(-cs\log(n/s))$, the following holds uniformly for all subsets $S\subset[n]$ with cardinality $|S|=s$, provided the sample size satisfies $m\geq Cs\log(n/s)$:
$\left\|\frac{1}{m}\sum_{i=1}^{m}\xi_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}-\mathbb{E}\left[\xi_{1}(\mathbf{a}_{1})_{S}(\mathbf{a}_{1})_{S}^{*}\right]\right\|$
$\leq$~ $C^{\prime}B\sqrt{\frac{s\log(n/s)}{m}},$ (A.7)
where $c,C,C^{\prime}$ are universal numerical constants.
Proof.
The proof follows the standard covering argument used for covariance matrix estimation (see, e.g., the proof of [23, Theorem 4.6.1]). The basic intuition is that the weighted bounded scalar will not essentially change the sub-exponential property of the random matrix. For simplicity, we denote
$\mathbf{E}_{i}:=\xi_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}-\mathbb{E}\left[\xi_{1}(\mathbf{a}_{1})_{S}(\mathbf{a}_{1})_{S}^{*}\right],$
and $\mathbf{E}:=\frac{1}{m}\sum_{i=1}^{m}\mathbf{E}_{i}$.
Step 1: Discretization via $\epsilon$-net. Fix a support $S$ with $|S|=s$. Let $\mathcal{N}$ be a $\frac{1}{4}$-net of the unit sphere in $\mathbb{C}^{|S|}$. The cardinality of the net is bounded by $|\mathcal{N}|\leq 9^{2|S|}\leq 9^{4s}$. From [23, Lemma 4.4.1], the operator norm of a Hermitian matrix $\mathbf{E}$ is controlled by the net
$\|\mathbf{E}\|\leq 2\sup_{u\in\mathcal{N}}|\mathbf{u}^{*}\mathbf{E}\mathbf{u}|.$ (A.8)
Thus, it suffices to bound the scalar deviation for each fixed $\mathbf{u}\in\mathcal{N}$.
Step 2: Scalar Concentration. Fix a unit vector $\mathbf{u}\in\mathcal{N}$. Consider the scalar random variable
$Z_{i}=\mathbf{u}^{*}\left(\xi_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right)\mathbf{u}=\xi_{i}|\langle(\mathbf{a}_{i})_{S},\mathbf{u}\rangle|^{2}.$
Since $(\mathbf{a}_{i})_{S}$ is standard Gaussian, the inner product $g=\langle(\mathbf{a}_{i})_{S},\mathbf{u}\rangle$ is a standard scalar complex Gaussian variable. The term $|g|^{2}$ follows an exponential distribution, which is sub-exponential with norm $\||g|^{2}\|_{\psi_{1}}\leq C_{sub}$. Because the weights are bounded ($|\xi_{i}|\leq B$), the variable $Z_{i}$ is sub-exponential with norm
$\|Z_{i}\|_{\psi_{1}}=\|\xi_{i}|g|^{2}\|_{\psi_{1}}\leq B\||g|^{2}\|_{\psi_{1}}\leq C_{sub}B.$
We now apply the scalar Bernstein inequality for sub-exponential variables [23, Theorem 2.8.1] and obtain
$\mathbb{P}\left(\left|\frac{1}{m}\sum_{i=1}^{m}(Z_{i}-\mathbb{E}Z_{i})\right|\geq t\right)$
$~~~~~~~~\leq 2\exp\left(-cm\min\left(\frac{t^{2}}{K^{2}},\frac{t}{K}\right)\right),$
where $K=\max_{i}\|Z_{i}\|_{\psi_{1}}\lesssim B$. Assuming large $m$ such that the sub-gaussian tail dominates, we set $t=C^{\prime}B\sqrt{\frac{s\log(n/s)}{m}}$ and get
$\mathbb{P}\Big(\Big|\underbrace{\frac{1}{m}\sum_{i=1}^{m}(Z_{i}-\mathbb{E}Z_{i})}_{\mathbf{u}^{*}\mathbf{E}_{S}\mathbf{u}}\Big|\geq t\Big)\leq~2\exp(-c^{\prime}s\log(n/s)).$
Note that the $B^{2}$ in the variance term cancels out, leaving the rate dependent on $m$ and $s$.
Step 3: Union Bound. We now control the probability that the operator norm exceeds $t$ for any valid support $S$. Taking the union bound over all $\binom{n}{s}$ supports and all $\mathbf{u}\in\mathcal{N}$ for each support yields
[…]
$\makebox[61.83356pt][c]{$\leq$}\binom{n}{s}\cdot 9^{4s}\cdot 2\exp\left(-c^{\prime}s\log(n/s)\right)$
$\makebox[61.83356pt][c]{$\leq$}\left(\frac{en}{s}\right)^{s}9^{4s}\cdot 2\exp\left(-c^{\prime}s\log(n/s)\right).$
By choosing the constant $C$ in the sample complexity $m\geq Cs\log(n/s)$ sufficiently large, the exponent $c^{\prime}s\log(n/s)$ dominates and thus ensures that the total probability is at most $2\exp(-cs\log(n/s))$. ∎
Based on the above lemmas, we proceed to show the similarity between the exponential spectral $\mathbf{Z}$ and its estimator $\mathbb{E}[\tilde{\mathbf{Z}}$]. First, we bound the operator norm of the difference matrix $\tilde{\mathbf{Z}}-\mathbb{E}[\tilde{\mathbf{Z}}$]. In fact, this is a direct application of Lemma 3.
@@
Let $\tilde{\mathbf{Z}}$ be defined as in (A.2) and its expectation is given by (A.3). For all $S\subseteq[n]$ with $|S|=s$, the following bound holds for the operator norm with probability exceeding $1-2\exp(-cs\log(n/s))$:
$\|\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\|\leq C\sqrt{\frac{s\log(n/s)}{m}}.$ (A.9)